OpenStudy (anonymous):

sin (theta) * cos (theta) = -.5

OpenStudy (anonymous):

\[\sin 2\theta=2\sin \theta \cos \theta =-1.0=-1=-\sin \frac{ \pi }{2 }=\sin \left(2n \pi-\frac{ \pi }{2 } \right),\],n=0,1,2,3 calculate further.